the area of a rectangle is 99 yd^2, and the length of the rectangle is 7 yd more than twice the width. what are the dimensions of the rectangle?
Answer by TimothyLamb(4379)

Lw = 99 sq.yd
L = 2w + 7
---
Lw = 99
(2w + 7)w = 99
2ww + 7w = 99
2ww + 7w - 99 = 0
---
the above quadratic equation is in standard form, with a=2, b=7, and c=-99
---
to solve the quadratic equation, plug this:
2 7 -99
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
the roots of the quadratic are:
-9
5.5
---
the negative root doesn't make sense as a length, so use the positive root:
---
w = 5.5
L = 18
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php